# Boaz Moerman

### v2

@@ 9.4. Geometric Campana and geometric Darmon points on Toric varieties
In [ShSt24], Shute and Streeter consider geometric Campana points and geometric Darmon points for Campana pairs $(X,D_{\mathbf{m}})$, where $X$ is a toric variety and $D_{\mathbf{m}}$ is torus-invariant. For such pairs, they prove an analogue of Manin’s conjecture on the number of geometric Campana points and geometric Darmon points of bounded log-anticanonical height. The invariants appearing in their asymptotic are direct analogues of those appearing in the conjecture [CLTT24, Conjecture 8.3] for Campana points. In particular, their result [ShSt24, Theorem 1.1] agrees with the prediction given in Conjecture 1.4.
@@ 9.5. Powerful values of norm forms
Now we will show that Conjecture 1.4 agrees with the results by Streeter [Str21, Theorem 1.1] on powerful values of norm forms. The other main result from that paper, [Str21, Theorem 1.4], is also compatible with Conjecture 1.4, as shown in [CLTT24, §8.5.2.].
@@
Under the assumption that $n$ is prime or $\gcd(n,m)=1$, Streeter obtained an asymptotic for the number of weak Campana points over $\mathcal{O}_{S}$ of bounded anticanonical height on the Campana pair $\left(X,\left(1-\frac{1}{m}\right)Z(N_{\omega})\right)$ with respect to the model $\mathcal{X}=\mathbb{P}^{n-1}_{\mathcal{O}_{S}}$, for any finite set of places $S$ containing an explicit set $S(\omega)\subset\Omega_{K}$. Here the adelic metrization $\mathcal{K}_{X}$ of $K_{X}$ used to define the height is the toric metric on the anticanonical divisor class as in [BaTs95, Theorem 2.1.6], where we view $X$ as a toric variety using the anisotropic torus $X\setminus Z(N_{\omega})$.
@@
For any integer $m\in\mathbb{Z}$, we compute the asymptotic formula predicted by Conjecture 1.4 and we show that it agrees with [Str21, Theorem 1.1] if $n$ is prime or $\gcd(n,m)=1$.
@@
Moreover, this prediction (with $Z=\emptyset$) agrees with [Str21, Theorem 1.1] if the assumptions of that theorem are satisfied.
@@
Now we will relate this with the results obtained by Streeter. Let $\mathcal{K}_{X}$ be the toric metrization of the canonical divisor as in [Str21, Definition 4.5.]. Under the additional assumption that $n$ and $m$ are coprime or $n$ is prime, [Str21, Theorem 1.1] states that
@@
The b-invariant: We first compare the exponent on the $\log B$ factor. In Streeter’s work [Str21, Proposition 5.8], he considers the set $G^{m}/S_{m}$, where $S_{m}$ acts by permuting the coordinates. The Galois group $G$ acts on $G^{m}/S_{m}$ by right multiplication of every element of a representative $m$-tuple. We choose an ordering on $G$. We have an isomorphism of $G$-sets
@@
Under this isomorphism, the $G$-subset $\Gamma_{M^{\circ}}$ corresponds to the $G$-subset of $G^{m}/S_{m}$ given by all $m$-tuples not containing some element of $G$. As in [Str21, Proposition 5.11], write $S^{\prime}(G,m)\subset(G^{m}/S_{m})/G$ for the quotient of this set by $G$. By the isomorphism, we see that $b(K,(X,M),-K_{X})=\#S^{\prime}(G,m)$. If we assume that $\gcd(m,n)=1$ or $n$ is prime as in [Str21, Theorem 1.1], then [Str21, Remark 5.12] gives $b(K,(X,M),-K_{X})=b(n,m)$. In particular, we have shown that the exponents on $B$ and $\log B$ in [Str21, Theorem 1.1] agree with our expectation.
@@
for all $b\in\operatorname{Br}_{1}(X,M^{\circ})/\operatorname{Br}(K)$, where $\tau_{X}$ is the Tamagawa measure on $X$ determined by $\mathcal{K}_{X}$. By [BaTs95, Proposition 3.4.4.], this is equal to
@@
Note that the factor $\frac{\operatorname{Res}_{s=1}\zeta_{K}(s)}{\operatorname{Res}_{s=1}\zeta_{E}(s)}$ appears here due to how the measure $\mu$ is normalized in [Str21, §3.2]. In order to prove that our constant $c$ agrees with the constant $\tilde{c}$, it only remains to relate our sum over Brauer elements with the sum over Hecke characters in [Str21].
@@
The torus $U$ is given as $U=R_{E/K}\mathbb{G}_{m}/\mathbb{G}_{m}$, where $R_{E/K}\mathbb{G}_{m}$ is the Weil restriction. Thus, an automorphic character on $U$ can be viewed as an automorphic character on $R_{E/K}\mathbb{G}_{m}$ whose restriction to $\mathbb{G}_{m}$ is trivial. In other words, we can view them as Hecke characters for $E$ whose restriction to $\mathbf{A}^{\times}_{K}$ is trivial. The group $\operatorname{Gal}(E/K)$ acts on the set of these Hecke characters by $\chi\mapsto\chi^{\sigma}$ for $\sigma\in\operatorname{Gal}(E/K)$ as in [Str21, Definition 2.23], where $\chi^{\sigma}(x)=\chi(\sigma(x))$ for all $x\in\mathbf{A}^{\times}_{E}$. We claim that $B=H[m]$ if $n=2$ and otherwise
@@
If $n>2$ is prime, then $\operatorname{Gal}(E/K)$ acts faithfully on the Galois modules $V_{\mathbf{w}}$ for $\mathbf{w}\in\Gamma_{M^{\circ}}$ so the map $V_{\mathbf{w}}\rightarrow\mathbb{Z}^{\operatorname{Gal}(E/K)}$ is an isomorphism. This implies that a character $\chi$ on $U$ satisfies $\chi_{\mathbf{w}}=1$ if and only if the character $\chi^{\mathbf{w}}$ obtained by the map $R_{E/K}\mathbb{G}_{m}\rightarrow U$ satisfies $\chi^{\mathbf{w}}=1$. This implies $B=B_{0}$. If $n=2$ or $n>2$ then $B$ includes the set of characters $\mathcal{U}_{0}$ in [Str21, Definition 5.19], which are the characters in $B$ which are invariant under the compact subgroup $\mathcal{K}$ as in [Str21, Definition 4.17]. Now [Str21, Lemma 5.3] implies that the only characters for which the integral does not vanish lie in $\mathcal{U}_{0}$, which shows that $c=\tilde{c}$.

### v3

@@ 9.4. Geometric Campana and geometric Darmon points on Toric varieties
In [45], Shute and Streeter consider geometric Campana points and geometric Darmon points for Campana pairs $(X,D_{\mathbf{m}})$, where $X$ is a toric variety and $D_{\mathbf{m}}$ is torus-invariant. For such pairs, they prove an analogue of Manin’s conjecture on the number of geometric Campana points and geometric Darmon points of bounded log-anticanonical height. The invariants appearing in their asymptotic are direct analogues of those appearing in the conjecture [13, Conjecture 8.3] for Campana points. In particular, their result [45, Theorem 1.1] agrees with the prediction given in Conjecture 1.4.
@@ 9.5. Powerful values of norm forms
Now we will show that Conjecture 1.4 agrees with the results by Streeter [49, Theorem 1.1] on powerful values of norm forms. The other main result from that paper, [49, Theorem 1.4], is also compatible with Conjecture 1.4, as shown in [13, §8.5.2.].
@@
Under the assumption that $n$ is prime or $\gcd(n,m)=1$, Streeter obtained an asymptotic for the number of weak Campana points over $\mathcal{O}_{S}$ of bounded anticanonical height on the Campana pair $\left(X,\left(1-\frac{1}{m}\right)Z(N_{\omega})\right)$ with respect to the model $\mathcal{X}=\mathbb{P}^{n-1}_{\mathcal{O}_{S}}$, for any finite set of places $S$ containing an explicit set $S(\omega)\subset\Omega_{K}$. Here the adelic metrization $\mathcal{K}_{X}$ of $K_{X}$ used to define the height is the toric metric on the anticanonical divisor class as in [4, Theorem 2.1.6], where we view $X$ as a toric variety using the anisotropic torus $X\setminus Z(N_{\omega})$.
@@
For any integer $m\in\mathbb{Z}$, we compute the asymptotic formula predicted by Conjecture 1.4 and we show that it agrees with [49, Theorem 1.1] if $n$ is prime or $\gcd(n,m)=1$.
@@
Moreover, this prediction (with $Z=\emptyset$) agrees with [49, Theorem 1.1] if the assumptions of that theorem are satisfied.
@@
Now we will relate this with the results obtained by Streeter. Let $\mathcal{K}_{X}$ be the toric metrization of the canonical divisor as in [49, Definition 4.5.]. Under the additional assumption that $n$ and $m$ are coprime or $n$ is prime, [49, Theorem 1.1] states that
@@
The b-invariant: We first compare the exponent on the $\log B$ factor. In Streeter’s work [49, Proposition 5.8], he considers the set $G^{m}/S_{m}$, where $S_{m}$ acts by permuting the coordinates. The Galois group $G$ acts on $G^{m}/S_{m}$ by right multiplication of every element of a representative $m$-tuple. We choose an ordering on $G$. We have an isomorphism of $G$-sets
@@
Under this isomorphism, the $G$-subset $\Gamma_{M^{\circ}}$ corresponds to the $G$-subset of $G^{m}/S_{m}$ given by all $m$-tuples not containing some element of $G$. As in [49, Proposition 5.11], write $S^{\prime}(G,m)\subset(G^{m}/S_{m})/G$ for the quotient of this set by $G$. By the isomorphism, we see that $b(K,(X,M),-K_{X})=\#S^{\prime}(G,m)$. If we assume that $\gcd(m,n)=1$ or $n$ is prime as in [49, Theorem 1.1], then [49, Remark 5.12] gives $b(K,(X,M),-K_{X})=b(n,m)$. In particular, we have shown that the exponents on $B$ and $\log B$ in [49, Theorem 1.1] agree with our expectation.
@@
for all $b\in\operatorname{Br}_{1}(X,M^{\circ})/\operatorname{Br}(K)$, where $\tau_{X}$ is the Tamagawa measure on $X$ determined by $\mathcal{K}_{X}$. By [4, Proposition 3.4.4.], this is equal to
@@
Note that the factor $\frac{\operatorname{Res}_{s=1}\zeta_{K}(s)}{\operatorname{Res}_{s=1}\zeta_{E}(s)}$ appears here due to how the measure $\mu$ is normalized in [49, §3.2]. In order to prove that our constant $c$ agrees with the constant $\tilde{c}$, it only remains to relate our sum over Brauer elements with the sum over Hecke characters in [49].
@@
The torus $U$ is given as $U=R_{E/K}\mathbb{G}_{m}/\mathbb{G}_{m}$, where $R_{E/K}\mathbb{G}_{m}$ is the Weil restriction. Thus, an automorphic character on $U$ can be viewed as an automorphic character on $R_{E/K}\mathbb{G}_{m}$ whose restriction to $\mathbb{G}_{m}$ is trivial. In other words, we can view them as Hecke characters for $E$ whose restriction to $\mathbf{A}^{\times}_{K}$ is trivial. The group $\operatorname{Gal}(E/K)$ acts on the set of these Hecke characters by $\chi\mapsto\chi^{\sigma}$ for $\sigma\in\operatorname{Gal}(E/K)$ as in [49, Definition 2.23], where $\chi^{\sigma}(x)=\chi(\sigma(x))$ for all $x\in\mathbf{A}^{\times}_{E}$. We claim that $B=H[m]$ if $n=2$ and otherwise
@@
If $n>2$ is prime, then $\operatorname{Gal}(E/K)$ acts faithfully on the Galois modules $V_{\mathbf{w}}$ for $\mathbf{w}\in\Gamma_{M^{\circ}}$ so the map $V_{\mathbf{w}}\rightarrow\mathbb{Z}^{\operatorname{Gal}(E/K)}$ is an isomorphism. This implies that a character $\chi$ on $U$ satisfies $\chi_{\mathbf{w}}=1$ if and only if the character $\chi^{\mathbf{w}}$ obtained by the map $R_{E/K}\mathbb{G}_{m}\rightarrow U$ satisfies $\chi^{\mathbf{w}}=1$. This implies $B=B_{0}$. If $n=2$ or $n>2$ then $B$ includes the set of characters $\mathcal{U}_{0}$ in [49, Definition 5.19], which are the characters in $B$ which are invariant under the compact subgroup $\mathcal{K}$ as in [49, Definition 4.17]. Now [49, Lemma 5.3] implies that the only characters for which the integral does not vanish lie in $\mathcal{U}_{0}$, which shows that $c=\tilde{c}$.